12) Let F(x,y) = yi + x2j. Evaluate Sc F. dr for the parabolic curve C: r(t) = ti + (4t-t?) Osts 3.
Answer #1

We have, have, F(x,y) = yi txs and c: 841 atît (4t- + ²) 7 , osta3 8(t)= i +14-et) and F(TH) = (44-4²) itten s Ed s F (54).